40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4993     Entry time: Tue Jul 19 23:39:11 2011
Author: Jamie, Jenne 
Type: Update 
Category: LSC 
Subject: Major overhaul of LSC rack; binary switching of whitening filters now working 

Yesterday we started going through the LSC binary whitening switching to make sure the new switching control in the LSC model was working.  Jenne and I hooked up a fancy home-brew white noise generator [0] into all of the LSC whitening filter inputs and started switching the whitening filters to see what would happen.  We found that some of the channels were switching, but the majority were not, or worse yet switching the wrong channels.  Upon closer inspection, and after finally finding the LSC wiring schematic, we found that the LSC rack cross-connect/back-plane cabling was pretty much a complete mess, and didn't at all correspond to the channel layout in Suresh's diagram.

Given that the back-plane wiring had to be almost completely redone, we decided to completely redo the LSC electronics layout, to be a little more self-consistent and to use the given space more efficiently.  We'll post an updated electronics drawing soon.  The LSC model was also updated to reflect the new layout.

We then went through and verified that all of the whitening switching was working with the new layout.  As described previously, the first filter in the PD input filter bank is used to control the switching.  We did indeed verify that all the switching is working, but we noticed that switching logic was inverted, such that the whitening filter engaged when the filter was turned off.  This was fixed in the model and all the switching logic was verified to be working as expected.

Everything has now been hooked back up, but we need to verify that we're getting all of the PD demodulated RF and DC outputs as expected.  We need to check the RF phases, as some of the RF cable lengths have changed.

[0] a 50k resistor

Links:

 

ELOG V3.1.3-